last updated June 29, 2015
Apache Point Observatory
New Mexico, USA
ICC and Software
ARCTIC Recent Technical Documents
Final Optics Report
ARCTIC Schedule and Progress Reports
ARCTIC Preliminary Design Review (PDR) Information
Scattered Light Studies for the ARC 3.5-m
Some useful information on baffling and stray light analysis. Fest Chapter 9 and Caldwell 1997.
June 27 -- FIRST LIGHT!! Optics look great, and all is well. Thanks to everyone for their hard work!!
June 27 -- Ready for commissioning!! Wiring is cleaned up and stable. ICC code is running and ready for exposures via telnet. Bias levels are tuned. Here we go!
June 25 -- ARCTIC is assembled and on the observing level. All looks well, with some finishing touches. Chip is currently running at 167K.
June 23 -- Installing the cable run. One of the optics does not fit in its clamp properly. New part installed tomorrow. Dewar is ready to be cooled after fixing swapped preamp boards. Temporary instrument cart is nearly ready.
June 21 -- Detector has been installed. All quadrants look good cold. Conor's software is working, and taking images in different configurations.Shutter code will be ready for pre-commissioning. Most of the critical machining is complete. Assembled on the obs level Wednesday.
June 3 -- Detector tested perfect. Discussing a new preamp board from Leach.
May 20 -- Detector at STA for testing.Tested fine warm, will test cold next week.
May 18 -- Lenses delivered, and meet our specs.
May 16 -- Shutter is installed in housing.
May 15 -- New instrument cart design upgrades. Will start ordering parts.
May 14 -- Problem with the UR quadrant of the detector. Did some troubleshooting with Leach.
May 11 -- Housing is anodized and aeroglazed. Optics will be shipped tomorrow.
April 30 -- Detector is cooling.
April 28 -- Detector is installed, along with the new heater
April 9 -- Our detector should be arriving today.
April 7 -- New SDSS filter set has arrived, and quality looks good.
April 7 -- Production shutter is working well, and will be placed in a freezer to test extreme conditions.
April 6 -- Housing has been taken to Albuquerque for anodizing. Expect it back in a few weeks.
April 3 -- STA got our extra tests done quickly, and the results look great.
April 2 -- STA sent a final detector report. There are some areas of concern, so we are asking for extra tests.
March 31 -- Cart design is nearly done. Open for discussions.
March 29 -- Since coming to APO, the CCD has been running too clod. An additional heater was ordered, and some modifications will be needed when we swap detectors.
March 27 -- Lenses are being coated. Expect them by the end of April if all goes to plan.
March 20 -- Housing is nearly done being machined. Will be ready for anodizing in about a week.
March 19 -- Filter wheel and shutter PCBs are complete.
March 2 -- ITL is finishing up thinning and coating our CCD. Should have final characterization soon.
February 25 -- SDSS filters expected to ship first week of April.
February 23 -- All machining is progressing well, ahead of schedule.
February 11 -- Dewar is pumped and cooled down in the ground level of the dome.
February 11 -- ARCTIC has arrived at APO. Condesation is gone, and there appears to be no damage to the CCD.
February 5 -- Camera is warming for transport to APO tomorrow.
February 5 -- Our CCD is at ITL for thinning and coating. Machining underway at CVE.
January 15 -- In-house shutter is being made, and off-the-shelf Bonn shutter has been ordered.
January 12 -- Order placed for the instrument housing.
December 8 -- Optics order has been placed, and manufacturing has started. Lab testing for the CCD has started.
November 17 -- Ion pump has been stable and instrument looks healthy, it is regulating itself properly.
November 6 -- Instrument arrives back at UW after CCD electronics integration at Leach.
October 30 -- CCD has been installed, Bill and Joe trained by Leach on how to tune it.
September 30 -- Dewar and associated equipment have been delivered to Leach's shop. Engineering CCD has been installed and is being cooled to see if there are any problems.
September 29 -- The ICC has been delivered to APO and is getting its software installed.
September 28 -- Thermal testing is complete, and the camera operated as expected. Without a heat load (like the CCD) and in a warm room, the temperatures came down to 150K and a vacuum of high 10e-7.
Expect CCD delivery beginning of October!
Dewar is in vacuum testing. Main goal was to check for leaks. Even with that, the final pressure at the vacuum pump was 3.5x10^-7 Torr, which is not bad.
Shutter prototesting is complete! Went well, except that a cylinder was damaged. It is being repaired.
November 5, 2013: Instrument Team adopts the name: ARC Telescope Imaging Camera or ARCTIC for the new instrument
Recent Telecon Notes